OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

A. x = 4, y = 11 B. x = -12, y = -5 C. x = -4, y = 3 D. x = 8, y = 3

OpenStudy (shamil98):

since y= -2x + 19 and y = x + 7 therefore: x + 7 = -2x + 19

OpenStudy (anonymous):

ok i dont understand

OpenStudy (shamil98):

y = -2x + 19 y = x+7 Y is both equal to x + 7 and -2x + 19 Substituting for Y x + 7 = -2x + 19

OpenStudy (anonymous):

oh ok what do i do after that

OpenStudy (shamil98):

Solve for x.

OpenStudy (anonymous):

i got it thanks

OpenStudy (shamil98):

no problem

OpenStudy (anonymous):

If you get x = 4, you are right. But your homework is about how to find it.

OpenStudy (anonymous):

Do not forget after to find y by substituting the value of x that you found in y = x+7. So if x = 4, you have y = x+7 = 4+7 = 11